OpenStudy (batman4ever):

the area of a rectangle is 16 inches. Determine the length and width if the length is 4 times the width.

OpenStudy (watchmath):

L=4W Area =L*W=4W^2=16 W^2=4 W=2 so L=4(2)=8.

OpenStudy (anonymous):

>.<

OpenStudy (batman4ever):

thank you!